Exercise 23.12 If the electric potential at point A is +30.0 V, what is the electric potential at point B Express your answer with the appropriate units. Constants An object with charge q-6.00109 Cis placed in a region of uniform electric field and is released from rest at point A. After the charge has moved to point B, 0.500 m to the right, it has kinetic energy 3.00 10-7 J VValue Units Submit Reguest Answer Part B What is the magnitude of the electric field? Express your answer with the appropriate units.
Answer #1

here,

charge , q = - 6* 10^-9 C

s = 0.5 m

a)

the potential at A , Va = 30 V

let the potential at B be Vb

using work energy theorm

work done by potential = kinetic energy gained

(Vb - Va ) * q = 3 * 10^-7

(Vb - 30) * 6 * 10^-9 = 3 * 10^-7

Vb = 80 V

the potential at B is 80 V

b)

the electric feild , E = (Vb - Va) /s

E = ( 80 - 30) /0.5

E = 100 N/C